OpenStudy (anonymous):

Simplify 8^-2 by making the exponent positive. A. 1/8 B. 1/8^2 C. 8^2 D. 64

OpenStudy (anonymous):

So when you have any number or function raised to a negative exponent, that is similarly the inverse of that number or function. For example: x^-1 = 1/x. 2^-1 = 1/2. Using this we can simply manipulate the 1 such that it carries. So. x^-2 = 1/(x^2), or 2^-2 = 1/(2^2). I believe the answer is obvious enough now.

OpenStudy (anonymous):